OpenStudy (anonymous):

Eliminate the parameter and write a rectangular equation for. x = t^2 + 3 y = 2t Just need some help on how to do this. Please show all of your work.

hartnn (hartnn):

just put t=y/2 in x=t^2+3

OpenStudy (anonymous):

Cant believe I didn't see it before, thanks.

OpenStudy (mathmate):

Eliminate the parameter t: \[ t=y/2\] \[x = \frac{y}{2}^2+3\]

hartnn (hartnn):

welcome :)